OpenStudy (anonymous):

find the vertex of the parabola (x+3)^2=4(y+2)

OpenStudy (anonymous):

Im sure the first part is just foiling out the x+3 which is x^2+6x+9

OpenStudy (anonymous):

don't "foil"

OpenStudy (anonymous):

bahaha ok, well, what should I do then?

OpenStudy (anonymous):

I just have to say your picture made me laugh so hard. lolol

OpenStudy (anonymous):

from the form you are in you see the first coordinate of the vertex is \(-3\) because that will make \((x+3)^2=0\) replace \(x\) by \(-3\) and see what you get for \(y\)

OpenStudy (anonymous):

me too, well played @satellite73. I need some laughs at this point. I'm getting burnt out on this stuff haha

OpenStudy (anonymous):

\[(x+3)^2=4(y+2)\] \[0=4(y+2)\iff y=-2\]

OpenStudy (anonymous):

it was displayed for you in the form you were given \[(x-h)^2=a(y-k)\] vertex is \((h,k)\)

OpenStudy (anonymous):

what became of the 4?

OpenStudy (anonymous):

oh, wait. nevermind. I guess it's just not needed.

OpenStudy (anonymous):

and as far as taking out the opposite as in the parenthesis, thats just to make it equal zero?